[Hallc_running] [New Logentry] Hall C update (Dec. 11)
gaskelld at jlab.org
gaskelld at jlab.org
Mon Dec 11 09:50:02 EST 2017
After the down time yesterday that began with work targeted at bringing up Hall A, beam returned in the evening (around 6:30 pm), and we began trying to optimize the SHMS tune.
The carbon 4.4 MeV state was not visible at 13.5 degrees, so the SHMS was rotated to 10 degrees. After some exploration of SHMS magnet settings, we have settled on a tune that has all 3 three quads set 5% higher than our initial starting tune. Right now, we are in the middle of a delta scan in the SHMS.
It was also found yesterday that the large Moller quads had been mis-set, resulting in a beam spot very large in x - after this was fixed, the beam spot became nice and round on the viewer just upstream of the target. I had hoped that this improved beam spot would result in harp scans that made more sense - unfortunately, I am still a little confused by the most recent harp scan results.
Dave Meekins also came in yesterday and fixed the issue with the target FSD preventing us from putting in the hydrogen target.
This afternoon (about 1 pm), we will take a Controlled Access for the following:
1. Steve Lassiter et al will work on HMS Q1 - this has the highest priority
2. EES may take a look at the harps at 3H07A and 3H07B
3. We will rotate the SHMS back to 7.5 degrees for measurement of the carbon elastic peak
There are a limited number of access keys, so any other work in the hall this afternoon is discouraged. Please coordinate with the Run Coordinator if you think you need to access the hall.
The ion chamber calibration issue from Saturday night is still unresolved, although our Ops liaison is coordinating with SSG to get this sorted out.
This is a plain text email for clients that cannot display HTML. The full logentry can be found online at https://logbooks.jlab.org/entry/3502495
-------------- next part --------------
An HTML attachment was scrubbed...
More information about the Hallc_running